Area of painting: A rectangular painting with a width of x centimeters has an area of x^2 + 50x square centimeters. Find a binomial that represents the length. Area = x^2 + 50x cm^2
Answer by stanbon(75887)

A rectangular painting with a width of x centimeters has an area of x^2 + 50x square centimeters. Find a binomial that represents the length. Area = x^2 + 50x cm^2
----------------------
Area = LW
x^2+50x = L(x)
Length = (x^2+50x)/x
Length = (x+50) cm
------
Cheers,
Stan H.